OpenStudy (lsdxmdmaddicthc):

1.8788^(1/2t) = 1.019 Can someone explain the steps involved?

Vocaloid (vocaloid):

take the log of each side 1.8788^(1/2t) = 1.019 log[1.8788^(1/2t)] = log(1.019) (1/2)t*log(1.8788) = log(1.019) t = 2*log(1.019)/log(1.8788) use a calculator to evalute t

Vocaloid (vocaloid):

what are you still confused about?

OpenStudy (lsdxmdmaddicthc):

I never learned how to dog log...I don't even know what it is...

OpenStudy (lsdxmdmaddicthc):

This is actually a Compound Interest Equation from Advanced Algebra with Financial Applications. Sorry I was AFK for a moment.

OpenStudy (lsdxmdmaddicthc):

A = P(1+r/n)^(nt) 9,000 = 4,790.23(1+0.038/2)^(2t) Solve for t

Vocaloid (vocaloid):

uh, I'm not sure how to explain logs in a simple way ;_; @dan815 help please ;_;